1.   Determine the magnetic field B in air at a distance of 5 cm of a long wire where a current of 8A flows. 

2.   Find the magnetic field produced by the solenoid if the number of loops is 500 and current passing through on it is 3A (Length of the solenoid is 40 cm). 



1
Expert's answer
2021-11-24T12:00:14-0500

1) We can find the magnetic field as follows:


"B=\dfrac{ \mu_0I}{2\pi r}=\dfrac{ 4\pi\times10^{-7}\ \dfrac{T\times m}{A}\times8\ A}{2\pi\times0.05\ m}=3.2\times10^{-5}\ T."

2) We can find the magnetic field produced by solenoid as follows:


"B=\dfrac{\mu_0NI}{L},""B=\dfrac{4\pi\times10^{-7}\ \dfrac{T\times m}{A}\times500\times3\ A}{0.4\ m}=4.7\times10^{-3}\ T."